# ProbTS: Benchmarking Point and Distributional Forecasting across Diverse Prediction Horizons

### v4

@@ 1 Introduction
Time-series forecasting has extensive applications in various industries, including traffic flow forecasting [42], renewable energy forecasting [65], and diverse forecasting demands in retail [9], finance [30], and climate [47]. It is crucial to provide forecasts across different prediction horizons, addressing both short- and long-term planning needs [14, 27, 5, 59]. Moreover, modern decision-making processes typically require not only point forecasts to quantify planning efficiency but also robust distributional estimations to manage uncertainty effectively [25, 31]. The fundamental need to produce accurate point and distributional forecasts across various horizons presents significant challenges to existing forecasting approaches.
@@
Despite the recent surge in building time-series foundation models over the past year [19, 54, 15, 11, 16, 26, 40, 21, 68, 3, 24], our understanding of their advantages and limitations, especially regarding essential forecasting needs like point and distributional forecasts across various horizons, is still limited. Many of these models claim to support arbitrary prediction horizons, employing different mechanisms that come with their own set of advantages and drawbacks. Among them, a select few offer capabilities for distributional forecasting, which, however, are typically confined to predefined closed-form distributions [54, 68] or discrete distributions with value quantization [3]. The emergence of these foundation models has brought about unprecedented zero-shot forecasting capabilities. Consequently, it is both timely and crucial to delve into an evaluation of their strengths and weaknesses, especially in relation to the fundamental forecasting needs mentioned earlier.
@@
Our research reveals that the specific data characteristics inherent to different forecasting requirements often play a crucial role in guiding the selection of model designs. As a result, it is crucial to have a comprehensive view of the essential forecasting needs. To aid in the analysis and interpretation of performance, we measure three essential data characteristics in ProbTS: the strength of trends and seasonality, and the complexity of the data distribution. Moreover, we have explicitly distinguished two fundamental methodological aspects within ProbTS that differentiate the existing forecasting models, largely influencing their pros and cons. The first aspect pertains to the approach for enabling distributional forecasts. This ranges from models focused solely on point forecasting [48, 39], to those employing pre-defined distribution heads based on certain assumptions about the data distribution [54, 68], and to sophisticated neural probabilistic modules designed to flexibly adapt to various data distributions [55, 60]. The second aspect is the decoding scheme used to generate multi-step forecasts, which can be either autoregressive (AR) or non-autoregressive (NAR).
By utilizing ProbTS, we conduct a systematic comparison between studies that focus on long-term point forecasting and those aimed at short-term distributional estimation, employing various forecasting horizons and evaluation metrics. Our overarching finding is that the strengths of these methods tend to diminish in scenarios they are rarely evaluated in, highlighting several important but unresolved research questions. Notably, while recent probabilistic forecasting approaches have shown proficiency in short-term distribution estimation, we discover that long-term distributional forecasting remains a significant challenge. This challenge stems from the difficulty in achieving adaptive distribution estimation that remains both efficient and effective as the prediction horizon extends—a topic that has not been thoroughly investigated in existing literature. Additionally, our analysis uncovers a clear divide in the choice of decoding schemes: most long-term point forecasting methods opt for NAR, whereas choices in short-term forecasting studies are more evenly split. Further investigation suggests that this preference may be linked to the challenge of error accumulation in AR decoding, particularly as the forecasting horizon lengthens or when dealing with strong trends. However, we also discover that AR decoding tends to perform better in scenarios with pronounced seasonality, indicating a potential area for further exploration in decoding strategies, especially for long-term forecasting. Together with the notable inefficiency of existing NAR-based probabilistic methods, such as CSDI [60], for long-term forecasts, our comparison between NAR and AR decoding encourages a wider exploration of decoding strategies in future research.
@@ 3 The ProbTS Tool
The model module in ProbTS explicitly differentiates critical methodological decisions, especially the decoding scheme (AR vs NAR) and the distributional estimation approach. Specifically, we employ the following mathematical formulation. We denote an element of a multivariate time series as , where represents the variate index and denotes the time index. At time step , we have a multivariate vector . Each is associated with covariates , which encapsulates auxiliary information about the observations. Given a length- forecast horizon, a length- observation history and corresponding covariates , the objective in time series forecasting is to generate the vector of future values . To represent point and distributional forecasting in a unified way, here we divide a model into an encoder and a forecaster . An encoder is tasked with generating expressive hidden states . Under autoregressive decoding scheme, encoder forecasts variates using their past values, which can be formulated as . Under the non-autoregressive scheme, the encoder generates all the forecasts in one step, which can be expressed as . A forecaster is employed either to directly estimate point forecasts as , or to perform sampling based on the estimated probabilistic distributions as .
@@ 4 Results and Analyses
[CRPS w.r.t. forecasting horizon.]
[CRPS w.r.t. trend strength.]
[CRPS w.r.t. seasonality strength.]
[The CRPS gap, calculated as , w.r.t. trend and seasonality strengths.]
@@ 4.1 Analyzing Classical Models for Time-series Forecasting
Existing probabilistic forecasting methods exhibit a balanced preference for both AR and NAR decoding schemes. For instance, TimeGrad employs an AR decoding scheme, whereas CSDI utilizes an NAR approach. This contrasts starkly with the aforementioned customized architectures, which solely opt for NAR decoding. These two types of decoding schemes, however, confront distinctive challenges when applied to long-term probabilistic forecasting. AR probabilistic models like TimeGrad tend to accumulate errors as they forecast further into the future or as the trend strength increases. This issue, which leads to a growing performance gap for longer forecasts or stronger trends, is illustrated in Figures 3 and 3. On the other hand, NAR probabilistic models, such as CSDI, face challenges related to memory usage and processing time in long-term forecasts, which are elaborated in Appendix Section D.6. Furthermore, as indicated in Table 10, even for relatively small-scale datasets, such as ETTm2 and ETTh1, CSDI’s performance in long-term scenarios is less than optimal. This suggests that its learning efficacy is also compromised by the extension of the forecasting horizon.
@@
Despite its drawbacks, AR decoding, as used in TimeGrad, excels in capturing strong seasonality, outperforming models like PatchTST in scenarios such as the Traffic dataset (Table 10). This advantage is further analyzed in Figures 3 and 3, showing AR’s increasing benefit with stronger seasonal patterns. This suggests AR’s potential in long-term forecasting could be revitalized with solutions to its error accumulation challenge in long horizons.
@@ 4.2 Analyzing Foundation Models for Universal Time-series Forecasting
Figure 4 illustrates how the performance of various time-series foundation models evolves in relation to expanding forecasting horizons. It is evident that for shorter horizons, AR-based foundation models such as TimesFM and Timer exhibit highly competitive performance, on par with NAR-based models like MOIRAI. However, the advantage of NAR-based decoding becomes increasingly apparent as the forecasting horizon lengthens, as demonstrated by the widening performance gap between TimesFM and MOIRAI. This trend is consistent with our earlier observation that AR-based methods could suffer from significant error accumulation when applied to long-term time-series forecasting. Given the inherent strengths of AR decoding, such as its superiority at capturing strong seasonality and its robust performance in certain short-term forecasting scenarios, it is clear that further research is warranted to explore ways to overcome its limitations in long-term forecasting contexts. This could potentially unlock new avenues for enhancing the versatility and effectiveness of AR-based time-series foundation models across a broader range of forecasting horizons.
@@ 5 Conclusion
The insights derived from our study open the door to numerous promising research directions. Addressing the shortcomings of AR and NAR decoding schemes, especially in their application across varying forecasting horizons, emerges as a critical area for future exploration. Innovating effective architecture designs that can navigate the intricacies of short-term forecasting challenges and devising efficient methods for long-term probabilistic forecasting stand out as urgent needs. For those time-series foundation models, developing mechanisms to combat error accumulation in AR-based methods could be very helpful. Equally important is the enhancement of models’ abilities to characterize complex data distributions, which could significantly improve the adaptability and effectiveness of foundation models. Beyond these technical endeavors, expanding the scope of datasets used for evaluation to encompass a wider range of real-world scenarios will be crucial for validating the robustness and versatility of future forecasting models. Lastly, integrating theoretical insights with empirical findings could provide a more holistic understanding of model behaviors, contributing to the development of more sophisticated and nuanced forecasting solutions.
Checklist
1.
For all authors…
Did you describe the limitations of your work? [Yes] See Sections 5.
2.
If you are including theoretical results…
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
4.
If your work uses existing assets, did you cite the creators? [Yes]
Did you mention the license of the assets? [Yes] See Appendix B.4.
5.
If you used crowdsourcing or conducted research with human subjects…
@@ A.3 A Comparison with Existing Libraries on the Coverage of Data, Model, and Metrics
ProbTS is a research toolkit designed to advance forecasting research across varied horizons, focusing on both point and distributional forecasting. To achieve these objectives, ProbTS includes state-of-the-art (SOTA) models, comprehensive evaluation protocols (point vs. distributional), and explores different methodological aspects of forecasting models, particularly in terms of distributional estimation methods and decoding schemes (AR vs. NAR). In Table 5, we provide a comprehensive comparison of ProbTS with existing libraries in terms of toolkit functionalities and the benchmarking aspects we aim to investigate.
@@ B.1.4 Quantifying Data Distribution Complexity
To differentiate between methods optimized for point or distributional forecasts, we aim to quantify the complexity of data distribution within a time-series segment whose window size equals the prediction horizon length. We propose that assessing non-Gaussianity, i.e., how closely the distribution of time-series values within that window resembles a Gaussian distribution, could serve as a meaningful measure. This is because point forecasting methods, optimized with mean squared loss, are essentially equivalent to probabilistic counterparts that include a Gaussian output head and employ maximum a posteriori estimation. This suggests that point forecasting methods inherently assume that time-series values adhere to a Gaussian distribution. In contrast, advanced probabilistic methods, which do not make prior assumptions about data distribution, can adapt to complex data distributions in a data-driven manner.
@@ Appendix D Additional Results and Experiments
D.1 Impact of Data Scale

### v5

@@ 1 Introduction
Time-series forecasting has extensive applications in various industries, including traffic flow forecasting [44], renewable energy forecasting [68], and diverse forecasting demands in retail [9], finance [30], physical system [40], and climate [49]. It is crucial to provide forecasts across different prediction horizons, addressing both short- and long-term planning needs [14, 27, 5, 62]. Moreover, modern decision-making processes typically require not only point forecasts to quantify planning efficiency but also robust distributional estimations to manage uncertainty effectively [25, 31]. The fundamental need to produce accurate point and distributional forecasts across various horizons presents significant challenges to existing forecasting approaches.
@@
Despite the recent surge in building time-series foundation models over the past year [19, 57, 15, 11, 16, 26, 42, 21, 71, 3, 24, 75], our understanding of their advantages and limitations, especially regarding essential forecasting needs like point and distributional forecasts across various horizons, is still limited. Many of these models claim to support arbitrary prediction horizons, employing different mechanisms that come with their own set of advantages and drawbacks. Among them, a select few offer capabilities for distributional forecasting, which, however, are typically confined to predefined closed-form distributions [57, 71] or discrete distributions with value quantization [3]. The emergence of these foundation models has brought about unprecedented zero-shot forecasting capabilities. Consequently, it is both timely and crucial to delve into an evaluation of their strengths and weaknesses, especially in relation to the fundamental forecasting needs mentioned earlier.
@@
Our research reveals that the specific data characteristics inherent to different forecasting requirements often play a crucial role in guiding the selection of model designs. As a result, it is crucial to have a comprehensive view of the essential forecasting needs. To aid in the analysis and interpretation of performance, we measure three essential data characteristics in ProbTS: the strength of trends and seasonality, and the complexity of the data distribution. Moreover, we have explicitly distinguished three fundamental methodological aspects within ProbTS that differentiate the existing forecasting models, largely influencing their pros and cons. The first aspect involves the approach to distributional forecasting, ranging from models focused on point forecasts [50, 41] to those using pre-defined distribution heads based on specific data assumptions [57, 71]. The second aspect is the decoding scheme used to generate multi-step forecasts, which can be either autoregressive (AR) or non-autoregressive (NAR). The third aspect pertains to the normalization choice, where the long-term point forecasting models typically employ reversible instance normalization (RevIN) [33] while short-term probabilistic ones often use mean scaling strategies [59, 58].
By utilizing ProbTS, we conduct a systematic comparison between studies that focus on long-term point forecasting and those aimed at short-term distributional estimation, employing various forecasting horizons and evaluation metrics. Our overarching finding is that the strengths of these methods tend to diminish in scenarios they are rarely evaluated in, highlighting several important but unresolved research questions. Notably, while recent probabilistic forecasting approaches have shown proficiency in short-term distribution estimation, we find that long-term distributional forecasting remains a significant challenge. This challenge stems from achieving distribution estimation that remains both efficient and effective as the prediction horizon extends—a topic that has not been thoroughly investigated in existing literature. Additionally, our analysis uncovers a clear divide in the choice of decoding schemes: most long-term point forecasting methods opt for NAR, whereas choices in short-term forecasting studies are more evenly split. Further investigation suggests that the preference for NAR methods stems from existing AR models’ difficulty in managing error accumulation, particularly over extended horizons with strong trends. However, we observe that a proper normalization strategy can significantly improve AR models in long-term forecasting, opening new possibilities for AR-based approaches. Moreover, AR decoding performs better in scenarios with pronounced seasonality, indicating potential for refining these strategies, particularly for long-term forecasts. Given the inefficiency of existing NAR-based probabilistic methods like CSDI, our comparison highlights the need for further exploration of decoding strategies in future research.
@@ 3 The ProbTS Tool
The model module in ProbTS explicitly differentiates critical methodological decisions, especially the decoding scheme (AR vs NAR) and the distributional estimation approach. Specifically, we employ the following mathematical formulation. We denote an element of a multivariate time series as , where represents the variate index and denotes the time index. At time step , we have a multivariate vector . Each is associated with covariates , which encapsulates auxiliary information about the observations. Given a length- forecast horizon, a length- observation history and corresponding covariates , the objective in time series forecasting is to generate the vector of future values . Based on established conventions, we categorize forecast as short-term if the horizon [58, 63], and long-term if [76, 50, 41], where represents the primary periodicity of the data (e.g., 24 for hourly frequency). To represent point and distributional forecasting in a unified way, here we divide a model into an encoder and a forecaster . An encoder is tasked with generating expressive hidden states . Under autoregressive decoding scheme, encoder forecasts variates using their past values: . Under the non-autoregressive scheme, the encoder generates all the forecasts in one step: . A forecaster is employed either to directly estimate point forecasts as , or to perform sampling based on the estimated probabilistic distributions as . In addition, the normalization choices utilized by different research branches vary, with a detailed analysis provided in Appendix D.1.
@@ 4 Results and Analyses
[CRPS w.r.t. forecast horizon.]
[CRPS of models (w/ ReVIN) and (w/ mean scaling) normalization.]
[CRPS w.r.t. trend and seasonality strength.]
[CRPS gap (TimeGrad.CRPS - CSDI.CRPS) w.r.t. trend and seasonality.]
Subplot (d) further investigates the combined effects of trend and seasonality, using lighter and smaller circles to indicate situations where AR-based models are favored over NAR-based ones.
@@ 4.1 Analyzing Classical Models for Time-series Forecasting
Existing probabilistic forecasting methods exhibit a balanced preference for both AR and NAR decoding schemes. For instance, TimeGrad employs an AR decoding scheme, whereas CSDI utilizes an NAR approach. This contrasts starkly with the aforementioned customized architectures, which solely opt for NAR decoding. These two types of decoding schemes, however, confront distinctive challenges when applied to long-term probabilistic forecasting. With original normalization strategy, AR probabilistic models like TimeGrad struggle with error accumulation, particularly as the forecast horizon extends or trends strengthen, the performance gap widens, as shown in Figures 3 and 3. On the other hand, NAR models such as CSDI encounter memory constraints in long-term forecasts (detailed in Appendix D.7). Moreover, Table 10 reveals that even on smaller datasets, such as ETTm2 and ETTh1, CSDI’s performance in long-term scenarios is less than optimal, indicating reduced learning efficiency by the extension of the forecasting horizon.
@@
Despite its drawbacks, AR-based models, as used in TimeGrad, excels in capturing strong seasonality, outperforming models like PatchTST in scenarios such as the Traffic dataset (Table 10). This advantage is further analyzed in Figures 3 and 3, showing AR’s increasing benefit with stronger seasonal patterns. This suggests AR’s potential in long-term forecasting could be revitalized with solutions to its error accumulation challenge in long horizons.
ReVIN’s Effectiveness in Long-term Forecasting with Exceptions.
RevIN significantly enhances AR-based models in long-term forecasting, as shown in Figures 3 and 7. Notably, on the ETTh1 dataset, GRU NVP (w/ RevIN) even outperforms PatchTST (w/ RevIN). While RevIN offers substantial improvements for most models across most datasets, it brings negative impact on the Traffic dataset. The Traffic dataset features strong seasonality but minimal trends, thus we speculate that the major distribution shift addressed by RevIN is related to normalizing the effect of trending. These findings indicate that normalizing the trending effect could be a direction to alleviate error accumulation of AR-based models in long-term forecasting. However, we also observe that RevIN does not seem to be an ideal match for the NAR probabilistic model. For instance, CSDI (w/ RevIN) performs worse than CSDI (w/ Scaling) on the Weather dataset. Further research in developing effective normalization strategies for NAR probabilistic models is necessary.
No Dominating Normalization Strategies in Short-term Forecasting.
While RevIN is effective for long-term scenarios, it does not adequately address the challenges faced by short-term probabilistic models. As shown in Table 14, RevIN fails to consistently deliver significant improvements for models such as CSDI, TimeGrad, and GRU NVP in short-term forecasting. The mean scaling strategy has proven to be the most reliable option for these models, explaining its widespread use. Although omitting instance-level normalization is occasionally acceptable, it can lead to significant issues, as seen with TimeGrad (without normalization) on the Wikipedia and Solar datasets, and GRU NVP (without normalization) on Electricity. We provide detailed analysis in Appendix D.1.
@@ 4.2 Analyzing Foundation Models for Universal Time-series Forecasting
Figure 4 illustrates how the performance of various time-series foundation models evolves in relation to expanding forecasting horizons. It is evident that for shorter horizons, AR-based foundation models such as TimesFM and Timer exhibit highly competitive performance, on par with NAR-based models like MOIRAI. However, the advantage of NAR-based decoding becomes increasingly apparent as the forecasting horizon lengthens, as demonstrated by the widening performance gap between TimesFM and MOIRAI. This trend is consistent with our earlier observation that without proper normalization strategies, AR-based methods could suffer from significant error accumulation when applied to long-term time-series forecasting. Given the inherent strengths of AR decoding, such as its superiority at capturing strong seasonality and its robust performance in certain short-term forecasting scenarios, it is clear that further research is warranted to explore ways to overcome its limitations in long-term forecasting contexts. This could potentially unlock new avenues for enhancing the versatility and effectiveness of AR-based time-series foundation models across a broader range of forecasting horizons.
@@ 5 Conclusion
The insights derived from our study open the door to numerous promising research directions. Addressing the shortcomings of AR and NAR decoding schemes, especially in their application across varying forecasting horizons, emerges as a critical area for future exploration. Innovating effective architecture designs that can navigate the intricacies of short-term forecasting challenges and devising efficient methods for long-term probabilistic forecasting stand out as urgent needs. For AR-based models, reducing error accumulation remains essential, with ReVIN-style normalization showing potential for improving long-term forecasting. Additionally, exploring effective normalization strategies for NAR-based probabilistic models is an underdeveloped yet promising area. Equally important is the enhancement of models’ abilities to characterize complex data distributions, which could significantly improve the adaptability and effectiveness of foundation models. Beyond these technical endeavors, expanding the scope of datasets used for evaluation to encompass a wider range of real-world scenarios will be crucial for validating the robustness and versatility of future forecasting models. Lastly, integrating theoretical insights with empirical findings could provide a more holistic understanding of model behaviors, contributing to the development of more sophisticated and nuanced forecasting solutions.
@@ A.3 A Comparison with Existing Libraries on the Coverage of Data, Model, and Metrics
ProbTS is a research toolkit designed to advance forecasting research across varied horizons, focusing on both point and distributional forecasting. To achieve these objectives, ProbTS includes state-of-the-art models, comprehensive evaluation protocols (point vs. distributional), and explores different methodological aspects of forecasting models, particularly in terms of distributional estimation methods and decoding schemes (AR vs. NAR). In Table 5, we provide a comprehensive comparison of ProbTS with existing libraries in terms of toolkit functionalities and the benchmarking aspects we aim to investigate.
@@ B.1.4 Quantifying Data Distribution Complexity
To differentiate between methods optimized for point or distributional forecasts, we aim to quantify the complexity of data distribution within a time-series segment whose window size equals the prediction horizon length. Such complexities may arise from the unpredictability of the data itself or from noises accidentally introduced during the data collection process [78].
We propose that assessing non-Gaussianity, i.e., how closely the distribution of time-series values within that window resembles a Gaussian distribution, could serve as a meaningful measure. This is because point forecasting methods, optimized with mean squared loss, are essentially equivalent to probabilistic counterparts that include a Gaussian output head and employ maximum a posteriori estimation. This suggests that point forecasting methods inherently assume that time-series values adhere to a Gaussian distribution. In contrast, advanced probabilistic methods, which do not make prior assumptions about data distribution, can adapt to complex data distributions in a data-driven manner.
@@ Appendix D Additional Results and Experiments
D.1 The Impact of Normalization
Normalization is a crucial aspect of time-series models, with different research branches adopting distinct strategies that can affect model performance across various data scenarios. Motivated by this, we provide a detailed analysis of different normalization methods in this section. Unless stated otherwise, our benchmarking follows the default normalization methods used by each model.
D.1.1 Different Normalization Choice Between Distinct Research Branches
In time series forecasting, normalization typically occurs in two stages. First, during preprocessing, dataset-level normalization is applied, where global statistics (e.g., mean and standard deviation) from the training set are used to normalize all time-series values. Then, a local normalization module might be used to perform instance-level normalization when feeding a batch of time-series segments into the model.
Different research branches prefer distinct instance-level normalization strategies. Long-term point forecasting models [50, 41] typically adopt the RevIN [33]. Given a batch of time-series segments within a lookback window, RevIN applies a per-series z-score normalization, augmented with learnable affine parameters. Its main advantage is its effectiveness in addressing distribution shifts, particularly in long-term forecasting.
In contrast, most short-term probabilistic forecasting models [58, 59] employ an ad-hoc but still effective normalization strategy. For example, given a batch of time-series segments (where is the number of variables and is the length of the lookback window), a per-series scaling is applied as to stabilize value ranges. For simplicity, we refer to this type of normalization as Mean Scaling.
We summarize the instance-level normalization choices originally used by each model in the Table 13.
Normalization Choice | Model |
ReVIN | iTransformer, PatchTST |
Mean Scaling | TimeGrad, GRU NVP |
w/o Norm | GRU, CSDI, DLinear |
Existing probabilistic models rarely use RevIN and are seldom combined with AR-based models that employ RevIN-style normalization. Similarly, the mean scaling strategy, commonly used in probabilistic forecasting models, is rarely applied to models designed for long-term forecasting. To better understand the effects of different normalization strategies, we selected representative models from both categories and combined them with three normalization methods: RevIN, Scaling (i.e., mean scaling), and w/o Norm (no instance-level normalization, using time-series values as provided by the dataset-level preprocessing). The results of these experiments are presented in Table 14, 15, 16, and 17.
Model | PatchTST | CSDI | TimeGrad | GRU NVP | ||||||||
ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | |
Electricity-S | - | |||||||||||
Exchange Rate-S | ||||||||||||
Solar-S | ||||||||||||
Traffic-S | ||||||||||||
Wikipedia-S |
Model | PatchTST | CSDI | TimeGrad | GRU NVP | ||||||||
ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | |
Electricity-S | - | |||||||||||
Exchange Rate-S | ||||||||||||
Solar-S | ||||||||||||
Traffic-S | ||||||||||||
Wikipedia-S |
Dataset | Pred | iTransformer | DLinear | PatchTST | CSDI | TimeGrad | GRU NVP | GRU | ||||||||||||||
len | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | |
ETTh1-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
ETTh2-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
ETTm1-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
ETTm2-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
Electricity-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
Exchange-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
ILI-L | 24 | |||||||||||||||||||||
36 | ||||||||||||||||||||||
48 | ||||||||||||||||||||||
60 | ||||||||||||||||||||||
Traffic-L | 96 | - | - | - | ||||||||||||||||||
192 | - | - | - | |||||||||||||||||||
336 | - | - | - | |||||||||||||||||||
720 | - | - | - | |||||||||||||||||||
Weather-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 |
Dataset | Pred | iTransformer | DLinear | PatchTST | CSDI | TimeGrad | GRU NVP | GRU | ||||||||||||||
len | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | w/o Norm | ReVIN | Scaling | |
ETTh1-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
ETTh2-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
ETTm1-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
ETTm2-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
Electricity-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
Exchange-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 | ||||||||||||||||||||||
ILI-L | 24 | |||||||||||||||||||||
36 | ||||||||||||||||||||||
48 | ||||||||||||||||||||||
60 | ||||||||||||||||||||||
Traffic-L | 96 | - | - | - | ||||||||||||||||||
192 | - | - | - | |||||||||||||||||||
336 | - | - | - | |||||||||||||||||||
720 | - | - | - | |||||||||||||||||||
Weather-L | 96 | |||||||||||||||||||||
192 | ||||||||||||||||||||||
336 | ||||||||||||||||||||||
720 |
D.1.2 Analysis of Instance-level Normalization
RevIN Significantly Improves Most Models in Long-term Forecasting Scenarios, with Some Exceptions. RevIN’s ability to mitigate the effects of data distribution shifts leads to significant performance improvements in most models for long-term forecasting, as shown in Table 16. This benefit extends beyond models like PatchTST and iTransformer, which originally employed RevIN, to others such as DLinear that do not inherently use this approach. Notably, RevIN has greatly enhanced AR-based models in long-term scenarios. For instance, on the ETT datasets, GRU NVP (w/ RevIN) outperforms even PatchTST (w/ RevIN), suggesting that normalizing trend effects can help reduce error accumulation in AR-based models.
However, RevIN can have a negative impact in certain cases. On the Traffic dataset, GRU (w/ ReVIN) and GRU NVP (w/ ReVIN) perform worse than without normalization, as shown in Table 16. Interestingly, this aligns with our analysis of data characteristics: the Traffic dataset displays strong seasonality but less trending. We speculate that RevIN’s effectiveness in other datasets stems from its ability to normalize trend-related distribution shifts, which is less relevant for the Traffic dataset. Additionally, RevIN appears less suited for NAR probabilistic models. For instance, CSDI (w/ RevIN) performs worse than CSDI (w/ Scaling) on the Weather, Electricity, Exchange, and ILI dataset. Further research is needed to develop more effective normalization strategies for NAR probabilistic models.
No Dominating Normalization Strategies in Short-term Forecasting. As shown in Table 14, RevIN does not consistently provide robust or significant improvements for models such as CSDI, TimeGrad, and GRU NVP in short-term forecasting. The Mean Scaling strategy, though empirical, proves to be the most reliable choice for these probabilistic models, likely explaining its widespread use. In some cases, instance-level normalization can be omitted, but this approach can lead to serious issues, as seen with TimeGrad (w/o normalization) on the Wikipedia and Solar datasets, and GRU NVP (w/o normalization) on the Electricity dataset. Developing effective instance-level normalization methods for complex data distributions in short-term forecasting remains an important yet often overlooked research direction.
[w/o Normalization.]
[ReVIN normalization.]
[Mean scaling normalization.]
[CRPS w.r.t. trend strength under each normalization choice.]
[CRPS w.r.t. seasonality strength under each normalization choice.]
[CRPS gap w.r.t. trend and seasonality strengths.]
[CRPS gap w.r.t. non-Gaussianity and seasonality strengths.]
[CRPS gap w.r.t. trend and seasonality strengths.]
[CRPS gap w.r.t. non-Gaussianity and seasonality strengths.]
D.2 The Impact of Data Scale